# FlowSearch: Advancing deep research with dynamic structured knowledge flow

### v1

@@ 3.2 Experiment Results
Table 1 and Figure 3 presents the performances of FlowSearch and its counterparts on GAIA, GPQA, HLE and TRQA. FlowSearch consistently achieves state-of-the-art results across all benchmarks, validating the effectiveness of the systematic design of FlowSearch.
3.2.1 General Question Answering
FlowSearch achieves state-of-the-art performance among agentic systems. On GAIA (Table 1), FlowSearch (o4-mini) outperforms both closed-source Manus (73.30%) and open-source OWL (69.70%), and shows strong robustness on Level 3 questions (50.00%). These results indicate that its iterative workflow combining knowledge planning, collection, and refinement is particularly effective for multi-hop and compositional reasoning. Its clear advantage over systems like OpenAI DR and MiroFlow further underscores the impact of structured and dynamic workflow design.
@@ 3.2.1 General Question Answering
Sequential Flow Planner Refiner Level 1 Level 2 Level 3 Avg Bio Chem Phys Avg
@@
90.56 76.74 50.00 76.96 84.21 79.57 96.51 87.37
Agentic systems consistently outperform pure LLMs on complex reasoning tasks. Larger base models like the Qwen series benefit from greater internal knowledge, but remain limited without structured reasoning. Even domain-finetuned models such as Intern-S1 lag behind agentic approaches. For instance, FlowSearch (o4-mini) achieves a GAIA score of 76.96%, far surpassing the same model without agency (16.97%), highlighting that structured task decomposition and flow-based execution are more critical than model size alone.
@@ 3.2.2 Multi-Disciplinary Research and Question Answering
General-purpose tools guided by Knowledge Flow can outperform specialized systems. On the HLE benchmark, FlowSearch (o4-mini) achieves the highest accuracy (30.80%), surpassing closed-source systems like OpenAI DR (26.60%) and Gemini Deep Research (26.90%). On TRQA, it reaches 77.9%, outperforming domain-specific agent Origene (60.1%) and scientific model Intern-S1 (49.4%). These results show that a well-structured general-purpose agent can effectively tackle complex scientific and cross-domain tasks.
@@ 3.4 Case Study and Visualization
Figure 4 illustrates the contrast between our knowledge-flow-based FlowSearch and the conventional sequential planning paradigm, represented by OWL [11], in addressing a scientific question. As shown in the figure, OWL decomposes the query into a linear sequence of subtasks—such as understanding, information collection, and identification—that are executed in order. While this pipeline is straightforward, it lacks mechanisms to preserve and integrate intermediate insights, which leads to the dilution of valuable evidence as the chain grows longer.
@@
4 Related Work
4.1 Agentic systems
Agentic systems with LLM have evolved from static prompting to perception–action loops, enabling systems to plan [24], act [5], and learn using external tools. Foundational approaches, such as interleaved reasoning–acting frameworks [5] and tree search planning [25], improve reliability in multi-step tasks, while reflective self-revision mechanisms [26] and external memory [27] enhance long-horizon consistency. Recent efforts like OpenHands [28] and OpenDevin [28] further expand agents’ action spaces and mitigate hallucinations through grounded APIs, tool calling, and software-execution platforms, and evaluate them on more realistic interactive benchmarks including AgentBoard [29], StuLife [30], and SWE-bench Verified [31]. Multi-agent orchestration involves role-specialized collaboration and negotiation [32, 33], replacing single-agent end-to-end optimization with a modular and scalable approach.
Despite these advances, most general-purpose agents target short to medium horizon tasks and interactive environments. Scientific research [21], however, requires handling long-horizon workflows, integrating diverse knowledge, and adapting strategies dynamically. This motivates the development of research-oriented agents, which focus on structured, adaptive, and knowledge-driven scientific inquiry.
4.2 Deep Research Agents
Recent advances in Deep Research (DR) agents extend LLMs from retrieval-augmented generation to dynamic, tool-driven research workflows. Early systems such as WebGPT [34] and Toolformer [7] explored web and API integration, while recent industrial solutions (e.g., OpenAI DR [21], Gemini DR [22], Grok DR [35], Perplexity DR [36]) incorporate adaptive planning, iterative retrieval, and multimodal reasoning. A key distinction is between static pipelines (e.g., AI Scientist [37], Agent Laboratory [38], and InternAgent [9]) and dynamic workflows, where plans evolve during execution. Dynamic workflows further divide into single-agent systems (Search-o1 [39], WebDancer [13], Qwen DeepResearcher [40]) that unify reasoning and tool use, and multi-agent systems (OpenManus [41], OWL [11], AWorld [42]) that distribute subtasks for parallel and specialised execution. While single-agent designs enable end-to-end optimization, multi-agent architectures offer modularity and scalability—crucial for complex research.
Recent studies also highlight graph-structured retrieval and adaptive workflows (e.g., GeAR [43], PANGU DeepDiver [44], Alita [33]), showing the benefit of explicit structures and self-evolving mechanisms for multi-hop reasoning. However, the existing DR agents still suffer from sequential bottlenecks and limited hierarchical decomposition, motivating frameworks like our FlowSearch that integrate multi-agent coordination with dynamic structured knowledge flow.
@@ 5 Conclusion
In this work, we introduce FlowSearch, a multi-agent deep research system built on a dynamic structured knowledge flow. By explicitly modeling dependencies among subproblems and key concepts, the system enables both deep reasoning within local regions of the knowledge flow and coherent knowledge propagation at a global level. The dynamic flow framework allows FlowSearch to iteratively plan, expand, and refine research workflows, supporting hierarchical task decomposition, parallel exploration, and adaptive strategy adjustment based on intermediate findings. These results highlight the effectiveness of combining structured knowledge flow planning with multi-agent orchestration, suggesting that such frameworks offer a promising direction for building autonomous, reflective, and scalable systems capable of tackling complex scientific research tasks.
@@ D.2 Report Generation
Multi-Agent AI Scientists in 2025
Introduction
With the rapid development of Large Language Models (LLMs) and Multi-Agent System (MAS) technologies, 2025 has seen the emergence of a series of “multi-agent AI scientist” platforms capable of collaboratively carrying out the entire process of scientific research. This report aims to provide a comprehensive reference for researchers and practitioners by synthesizing the latest definitions, historical background, core methodologies, datasets and evaluation benchmarks, real-world application cases, as well as the challenges faced and directions for future development in the field of multi-agent AI scientists in 2025.
Overview of Methods / Workflows
This report is based on a comprehensive analysis of multi-node workflows, drawing mainly on three core sections: (i) background and methods review, (ii) datasets and applications summary, and (iii) analysis of challenges and future directions. Each section compiles key insights from academic papers, industry blogs, white papers, technical reports, and authoritative surveys, preserving original references and data rigorously.
Findings and Synthesis
1. Background and Methods
1.1 Definitions
  • Multi-Agent System (MAS) — a computerized system composed of multiple interacting agents (software, robots, humans, or hybrids). Core features include autonomy, local views (no global state), and decentralization. Typical research themes encompass agent communication and coordination, distributed problem solving, multi-agent learning, DCOPs, BDI architectures, and LLM-based MAS.
  • Multi-Agent “AI Scientist” — a specialized MAS in which agents collaborate to complete the end-to-end scientific research process: hypothesis generation, experimental design, data interpretation, manuscript writing, and peer review. These systems are built upon advances in LLMs: each agent is usually an LLM or LLM+tools, specialized for a specific subtask of the scientific method.
1.2 Historical Development and Background
(a) Milestones in Laboratory Automation
  • 1875: First reports of automated scientific equipment, custom-built by scientists to address lab problems.
  • Post–WWII: Commercial vendors began offering increasingly complex automated lab equipment.
  • 1981–1983: Dr. Masahide Sasaki established the first fully automated laboratory.
  • 1993: Dr. Rod Markin developed the first clinical laboratory management system and led CTASSC (Clinical Testing Automation Standards Steering Committee).
  • 2004: The NIH Roadmap emphasized molecular libraries and imaging technologies, driving large-scale automation in biomedical research.
  • Mid-2010s: Rise of commercial “on-demand” remote labs (e.g., Emerald Cloud Lab, Strateos); studies showed that over 90% of methods in biomedical papers could be accessed through such services.
  • Low-cost automation era: Scripting languages (e.g., AutoIt) and open-source modules (LEGO, 3D printers) lowered the barrier for small labs.
(b) Early “Robot Scientist” Prototypes
  • Adam (Robot Scientist) by Ross King et al. — the first system to autonomously discover new knowledge in yeast functional genomics. Capabilities included hypothesis generation, experimental design, robotic execution, results interpretation, and iterative experimentation (domain: Saccharomyces cerevisiae).
  • Eve — a successor system focused on automated drug screening and reproducibility testing in cancer research.
(c) Conceptual Basis: MAS
Traditional MAS research emphasizes communication, coordination, distributed problem solving, multi-agent learning, DCOPs, and BDI-style architectures.
(d) Shift to AI-Driven MAS for Scientific Research
Advances in LLMs have enabled specialized agents to handle scientific subtasks (literature review, hypothesis generation, coding, data analysis, writing, peer review). Pipelines evolved from single tools to multi-agent workflows orchestrated by a central controller, simulating the iterative cycle of the scientific method. Human intervention is minimized: humans primarily handle wet-lab work and final oversight, while AI agents manage creativity, design, execution, analysis, and writing.
(e) Representative Prototypes and Frameworks before 2025
  • “Lab in the Loop” (FutureHouse & Oxford) — System: Robin. Agents: Crow & Falcon (literature analysis; Crow summarizes topics and proposes experimental designs; Falcon writes extended technical reports), Finch (data analysis of raw data such as RNA-seq and flow cytometry), and a Tournament Judge (hypothesis ranking using Bradley–Terry paired comparisons). Closed-loop iterations proceed as Crow/Falcon $\rightarrow$ Finch $\rightarrow$ Judge. In roughly ten weeks, the system completed drug repurposing studies for ten diseases and identified candidate compounds for dry age-related macular degeneration (dry AMD). Humans performed wet-lab experiments and finalized the manuscript.
  • Zochi (Intology) — a pipeline of retrieval, hypothesis generation, code execution, analysis, writing, and optional automated peer review. Human intervention occurs only ahead of major computation and during final polishing (figures, citations, formatting). Reported outcomes include ACL 2025 acceptance (average reviewer score 4.0), and projects such as CS-ReFT and Tempest/Siege.
  • The AI Scientist (Sakana.ai) — a four-stage loop: Idea Generation (guided by code templates and Semantic Scholar search), Experimental Iteration (running and logging code experiments with quantitative/visual outputs), Paper Write-Up (LaTeX drafts with auto-citations), and Automated Peer Review (LLM-driven near-human feedback). The loop is open-ended (each round is archived and informs the next). Estimated cost is approximately \$15 per paper. Known limitations include lack of precise layout/typography control, occasional implementation bugs, and security concerns when running unsandboxed code.
1.3 Common Architectural Themes
  • Agent specialization across subtasks: literature, hypothesis, coding, analysis, writing, and review.
  • Workflow orchestration: centralized coordination following the scientific method’s sequence.
  • Iterative cycles: hypothesis $\rightarrow$ experiment $\rightarrow$ analysis $\rightarrow$ refinement.
  • Minimal human intervention: humans focus on wet-lab work and final polishing.
  • Automated quality control: hypothesis ranking (e.g., Bradley–Terry), LLM-based peer review, and meta-analysis.
1.4 Evolving Standards and Frameworks
  • Legacy MAS protocols (e.g., KQML, ACL, FIPA) are increasingly complemented or replaced by LLM-driven frameworks (e.g., CAMEL).
  • Safety and ethics considerations include sandboxing, traceability, transparent AI attribution, and biosafety/software-security safeguards.
1.5 Pioneering Works
  1. 1.
    Lu et al., “The AI Scientist: Towards Fully Automated Open-Ended Scientific Discovery” (arXiv:2408.06292, Sep 2024). First comprehensive LLM-agent framework spanning idea generation, code execution, visualization, manuscript writing, and review loops; estimated cost <<\$15 per paper; code: https://github.com/SakanaAI/AI-Scientist.
  2. 2.
    Su et al., “Many Heads Are Better Than One: Improved Scientific Idea Generation by an LLM-Based Multi-Agent System” (arXiv:2410.09403, May 2025; ACL 2025). Introduces Virtual Scientists (VirSci)—a set of specialized LLMs that generate, evaluate, and optimize research ideas—outperforming single-agent baselines.
  3. 3.
    Ghareeb et al., “Robin: A Multi-Agent System for Automating Scientific Discovery” (arXiv:2505.13400, May 2025 submission). A closed-loop “lab-in-the-loop” MAS integrating literature review, hypothesis generation, experiment design, data analysis, and iterative refinement.
1.6 Representative Methods and Model Architectures
(a) System Goals and Benefits
These systems address dynamic, dependency-rich research tasks; enable parallel exploration; compress and merge context; and, compared with single-agent baselines, can deliver performance gains on the order of $\sim$90.2% at the cost of roughly $15\times$ token usage.
(b) Orchestrator–Worker Pattern
A Lead (Orchestrator) Agent analyzes user queries, drafts strategies, persists plans, spawns/monitors subagents, synthesizes outputs, and manages loop termination. Subagents (Workers) execute specialized tasks in parallel (e.g., web retrieval, data extraction) using interleaved thinking strategies to optimize queries and return structured results. A dedicated CitationAgent can post-process documents to produce precise citations.
(c) Model Backends and Tool Interfaces
Representative configurations include a powerful model for the lead agent (e.g., Claude Opus 4) and more efficient models for subagents (e.g., Claude Sonnet 4). Tool access is mediated by the Model Context Protocol (MCP) for network search, document retrieval, and workspace APIs. Dynamic multi-step retrieval often replaces static RAG, guided by in-prompt tool-selection heuristics.
(d) Prompt-Engineering Principles
  1. 1.
    Simulate agent execution to expose prompt failure modes.
  2. 2.
    Teach the orchestrator to delegate: clarify goals, output schemas, and tool guidance.
  3. 3.
    Scale the number of agents and tool calls with query complexity.
  4. 4.
    Design clean, purpose-built tools.
  5. 5.
    Let agents diagnose and improve other agents’ prompts and tool use.
  6. 6.
    Start broad, then progressively focus query strategies.
  7. 7.
    Use extended thinking as an explicit scratchpad.
  8. 8.
    Parallelize tool calls to reduce latency by up to 90%.
(e) Evaluation Strategies
Small-sample trials ($\sim$20 queries) assess prompt and tool efficacy; LLM-as-judge scoring covers factuality, citation accuracy, completeness, source quality, and efficiency. Human evaluation supplements these metrics to surface hallucinations and biases. Outcome-centric evaluations focus on persistent state changes.
(f) Production Engineering and Reliability
Stateful execution, checkpointing and retries; full tracing of agent decisions and tool calls; progressive (“rainbow”) deployment; external storage for plan digests and artifacts; and artifact systems to reduce token costs.
(g) Architectural Challenges and Future Directions
Synchronizing subagents remains a bottleneck (driving interest in asynchronous coordination). High operational costs encourage focusing on high-value domains. Coupling across subtasks and global-context management also remain open challenges.
2. Datasets, Benchmarks, and Application Cases
2.1 Datasets
(a) AI Idea Bench 2025 (arXiv:2504.14191)
A large-scale benchmark for idea generation covering recent venues (ICLR 2025, CVPR 2024, ECCV 2024, NeurIPS 2024, ICML 2024, NAACL 2024, EMNLP 2024, ACL 2024). It aggregates 3,495 target AI papers (post–Oct 10, 2023), plus the five most-cited works for each. Tasks include Idea Multiple-Choice Evaluation (IMCQ), Idea-to-Idea (I2I), and Idea-to-Topic (I2T). Representative results are summarized in Table 7.
Table 7: Representative AI Idea Bench 2025 results (illustrative extraction).
Method I2T IMCQ (A/B) Novelty / Feasibility Notes
AI-Scientist 5.0/5.0 3.591 / 2.734 $17.003\times 10^{-3}$ (highest novelty)
AI-Researcher 4.994/4.995 2.807 / 2.024 $9.728\times 10^{-3}$ (overall feasibility)
VIRSCI 4.974/4.983 2.937 / 2.123 strongest protocol alignment
SCIPIP 4.986/— 2.437 / —
(b) Open-Source Image Datasets (DatasetAgent; arXiv:2507.08648)
A multi-agent pipeline where four specialized agents use multimodal LLMs and an image-optimization toolkit to expand or build datasets according to user specifications. Downstream uses include image classification, detection, and segmentation. Specific source datasets (e.g., COCO, ImageNet) are not enumerated here.
2.2 Benchmarks and Evaluation Indicators
  • Stanford HAI 2025 AI Index Report: a general AI benchmark compendium including emerging multi-agent indicators. Highlights include improvements on MMMU (+18.8pp), GPQA (+48.9pp), and SWE-bench (+67.3pp). Responsibility-focused benchmarks covered include HELM Safety, AIR-Bench, and FACTS. The performance gap between open and closed weights narrows from 8% to 1.7%. Multi-agent findings note that language-model agents can surpass humans on time-limited programming tasks.
  • DatasetAgent (arXiv:2507.08648): evaluated via downstream CV tasks (classification, detection, segmentation); the abstract reports methodology without detailed aggregate metrics.
2.3 Real-World Applications and Case Studies
(a) Google AI Co-Scientist
A Gemini 2.0–based coalition of agents (Generation, Reflection, Ranking, Evolution, Proximity, Meta-review) supervised by a Supervisor agent. Capabilities include iterative self-debate and tournament-style ranking (Elo-like), automated literature review, hypothesis generation, proposal writing, experimental design, and recursive self-critique. Demonstrations:
  • AML drug repurposing: AI proposed KIRA6; validated in vitro across multiple AML cell lines, showing significant tumor-suppression at clinically relevant concentrations.
  • Liver fibrosis targets: identified epigenetic targets with anti-fibrotic activity in human liver organoids (collaboration with Stanford; $p<0.01$).
  • Antimicrobial resistance mechanisms: reproduced cf-PICI host-range expansion mechanisms with timelines consistent with unpublished experimental results (Fleming Initiative & Imperial College London).
(b) Stanford Virtual Lab of AI Scientists
An AI Principal Investigator agent coordinates subagents spanning immunology, computational biology, and machine learning, along with critic agents. Workflow: a human states a scientific challenge, the AI PI assigns roles, the agents conduct a parallel “lab meeting” lasting seconds to minutes and record meeting minutes, and tools such as AlphaFold assist in experimental design. A case study on SARS-CoV-2 vaccine design selected nanobody candidates that were validated for stability, variant binding, and low off-target effects by Chan Zuckerberg Biohub, demonstrating the acceleration provided by autonomous, high-throughput virtual experimentation.
(c) PriM: Principle-Inspired Material Discovery (arXiv:2504.08810)
A language-reasoning multi-agent system combining domain principles to guide exploration of chemical space via hypothesis generation and round-table MAS discussion. Demonstrated higher discovery rates and improved material properties in a nano-helix materials case compared with baselines. Code/data: https://github.com/amair-lab/PriM.
(d) SparksMatter: Autonomous Inorganic Materials Discovery (arXiv:2508.02956)
[…]
  • Green thermoelectrics: proposed Zintl phase CaMg2Si2 with energy above the convex hull $\leq 0.05\,\mathrm{eV/atom}$; predicted band gap $0.556\,\mathrm{eV}$ and bulk modulus $54.5\,\mathrm{GPa}$, with follow-up validation plans.
  • Soft inorganic semiconductors: identified Hg2MgRb2 (bulk modulus $K=19.94\,\mathrm{GPa}$; band gap $1.52\,\mathrm{eV}$; energy above hull $0.036\,\mathrm{eV/atom}$) including structure and synthesis routes.
  • Lead-free perovskite oxides: selected KNaNb2O6 isotypes ( $E_{\mathrm{hull}}<0.03\,\mathrm{eV/atom}$; band gap $\approx 2.4\,\mathrm{eV}$; bulk modulus $\approx 98\,\mathrm{GPa}$ ) with ferroelectric potential and validation strategies.
Benchmarking indicates superior novelty, depth, and rigor relative to OpenAI o3 / o3-deep-research / o4-mini-deep-research baselines; the authors identify scientific limitations and propose concrete improvements. Code/data: https://github.com/lamm-mit/SparksMatter.
Challenges, Limitations, and Future Work
3.1 Major Challenges and Systemic Risks
(IONI.ai Compliance Blog, Feb 15, 2025)
  1. 1.
    Agent failures: errors from a single agent can cascade. Mitigation: data-governance policies, pre-deployment testing, and fault isolation.
  2. 2.
    Coordination complexity: task allocation and messaging protocols are hard, especially with diverse agent roles. Mitigation: layered, sequential, or bidirectional coordination frameworks.
  3. 3.
    Unpredictable / emergent behaviors: decentralized agents may conflict or act unexpectedly. Mitigation: real-time monitoring, conflict-resolution protocols, and human-in-the-loop control.
  4. 4.
    Development and deployment challenges: deep planning is required for multi-agent integration. Best practices: phased rollouts, continuous monitoring, and simulation-based validation.
  5. 5.
    Scalability issues: increasing agent counts and task complexity introduce compute/network bottlenecks. Solutions: dynamic resource allocation, infrastructure optimization, scalable coordination algorithms.
  6. 6.
    Security and privacy: sensitive data traversing multiple agents increases risk. Measures: end-to-end encryption, strict access controls, and periodic security audits.
  7. 7.
    Troubleshooting complexity: emergent behaviors and dependencies complicate debugging. Tools: comprehensive logging, distributed tracing, unified monitoring dashboards.
  8. 8.
    Ethics and bias: data biases can be amplified in hypothesis generation and experimental design. Mitigation: bias audits, fairness-aware algorithms, and diverse training corpora.
3.2 Domain-Specific Limitations in Science
(Google Research Blog, Feb 19, 2025)
  1. 1.
    Literature coverage: cross-disciplinary or non-indexed publications may be missed. Remedy: widen retrieval modules to broader scientific corpora.
  2. 2.
    Fact-checking: proposals and hypotheses may contain factual errors. Remedy: integrate external verification tools and authoritative databases.
  3. 3.
    Tool interoperability: reliance on built-in tools limits domain-specific simulations. Remedy: API-driven plugins for scientific software.
  4. 4.
    Evaluation misalignment: Elo-like metrics can diverge from expert judgment. Remedy: hybrid evaluation combining automated scoring with domain-expert feedback.
  5. 5.
    Limited expert validation: early studies involved small expert cohorts and narrow targets. Remedy: larger-scale, cross-disciplinary expert evaluations to enhance generalization.
3.3 Future Work and Research Directions
  1. 1.
    Modular collaboration architectures: decompose the scientific pipeline into Generation, Reflection, Ranking, Evolution, Proximity, and Meta-review modules with encapsulation, transparency, and adaptivity; support tight human–AI collaboration.
  2. 2.
    Fusion of generative AI and MAS: integrate LLM components safely, strengthen coordination and distributed decision-making, ensure ethics and traceability, and optimize resource efficiency.
  3. 3.
    Cognitive architectures and distributed MAS: explore edge computing and network heterogeneity to improve robustness; investigate architectures such as Soar and ACT-R for commonsense reasoning and long-horizon goal management.
  4. 4.
    External tools and fact verification: expand cross-disciplinary retrieval; introduce fact-checking engines and authoritative knowledge bases; build plugin interfaces for scientific software.
  5. 5.
    Evaluation and expert collaboration: move beyond Elo-style metrics with hybrid frameworks; organize broader, more diverse expert reviews; implement Trusted Tester programs.
  6. 6.
    Governance and community standards: develop interoperability protocols, ethics and compliance norms, and reproducibility standards for multi-agent scientist systems; clarify legal accountability and funding mechanisms.
Conclusion
By 2025, multi-agent AI scientist technologies have made significant progress across definitions and methods, representative systems, datasets and evaluation benchmarks, and real-world applications, while also clarifying key challenges and avenues for future research. Shared characteristics include agent specialization, pipeline-style workflow orchestration, closed-loop iteration, and minimal human intervention. Real-world cases spanning drug repurposing, vaccine design, and materials discovery demonstrate the potential of multi-agent systems to accelerate scientific discovery. Future work should focus on robustness, safety, ethical compliance, and scalable evaluation frameworks to enable sustainable transition from prototypes to large-scale deployment.

### v2

@@ 4.2 Experiment Results
Table 1 and Fig. 3 present the performance of FlowSearch and its counterparts on GAIA, GPQA, HLE, and TRQA. FlowSearch consistently achieves competitive results across all benchmarks without training an additional base model, validating the effectiveness of its systematic design.
4.2.1 Agentic Benchmark
FlowSearch achieves state-of-the-art performance among agentic systems. On GAIA (Table 1), FlowSearch (o4-mini) outperforms both closed-source Manus (73.30%) and leading open-source agentic models Mirothinker (81.9%) and Tongyi-DR (70.9%), even though they are specifically trained and evaluated only on the GAIA text-only subset. FlowSearch also shows strong robustness on Level 3 questions (61.54%). These results indicate that its iterative workflow combining knowledge planning, collection, and refinement is particularly effective for multi-hop and compositional reasoning. Its clear advantage over systems like OpenAI DR and MiroFlow further underscores the impact of structured and dynamic workflow design.
@@ 4.2.1 Agentic Benchmark
Sequential Planner Flow Planner Refiner Level 1 Level 2 Level 3 Avg Bio Chem Phys Avg
@@
92.45 82.56 61.54 82.42 84.21 79.57 96.51 87.37
Agentic systems consistently outperform pure LLMs on complex reasoning tasks. Larger base models like the Qwen series benefit from greater internal knowledge, but remain limited without structured reasoning. Even models fine-tuned for scientific reasoning, such as Intern-S1, lag behind agentic approaches. For instance, FlowSearch with o4-mini achieves a score of 82.42% on GAIA, far surpassing the same model (o4-mini) without agency (16.97%), highlighting that structured task decomposition and flow-based execution are more critical than model size alone.
@@ 4.2.2 Multi-Disciplinary Research and Question Answering
Although scientific benchmarks are generally knowledge-intensive and favor LLMs, FlowSearch still achieves competitive performance.
@@
General-purpose tools guided by Knowledge Flow can outperform specialized systems. On the HLE benchmark, FlowSearch (o4-mini) achieves the highest accuracy among training-free methods at 34.52%, demonstrating performance comparable to leading trained agentic models such as MiroThinker (37.7%) and Tongyi-DR (32.9%) on the text-only subset, as well as closed-source systems including OpenAI DR (26.60%) and Gemini Deep Research (26.90%). On TRQA, FlowSearch reaches 77.9%, outperforming domain-specific agent Origene (60.1%) and scientific multi-modal model Intern-S1 (49.4%). These results show that a well-structured general-purpose agent system can effectively tackle complicated scientific and cross-domain tasks.
@@ 4.4 Case Study and Visualization
Fig. 4 illustrates the contrast between our knowledge-flow-based FlowSearch and the conventional sequential planning paradigm, represented by OWL [undefj], in addressing a scientific question. As shown in the figure, OWL decomposes the query into a linear sequence of subtasks—such as understanding, information collection, and identification—that are executed in order. While this pipeline is straightforward, it lacks mechanisms to preserve and integrate intermediate insights, which leads to the dilution of valuable evidence as the chain grows longer.
@@ 5 Conclusion
In this work, we have presented FlowSearch, a multi-agent deep research system built on a dynamic structured knowledge flow. By explicitly modeling dependencies among subproblems and key concepts, the system enables both deep reasoning within local regions of the knowledge flow and coherent knowledge propagation at a global level. The dynamic flow framework allows FlowSearch to iteratively plan, expand, and refine research workflows, supporting hierarchical task decomposition, parallel exploration, and adaptive strategy adjustment based on intermediate findings. These results highlight the effectiveness of combining structured knowledge flow planning with multi-agent orchestration, suggesting that such frameworks offer a promising direction for building autonomous, reflective, and scalable systems capable of tackling complex scientific research tasks.